Skip to main content

AAVSOnet status, 2017-03-26

1 post / 0 new
HQA
HQA's picture
AAVSOnet status, 2017-03-26

Here is the status of the various systems.

BSM_Berry  operational

BSM_Hamren operational

BSM_HQ operational

BSM_NM operational, with some caveats.  There appears to be a camera moisture problem; Bill Stein hopes to get up this week and replace the dessicant.  There were ~16 test nights during November/December that will be sent to users in the next few days.  These nights used local calibration (dark subtraction/flatfielding) rather than calibration at HQ, and so there may be a few differences.  There was also a problem on 170201, 170202, 170204, and 170205 with an incorrect bias setting.  These nights will be reprocessed in a few days.  We recommend not doing any analysis on those 4 nights until the new processing is available.

BSM_S operational

SRO  currently down.  A circuit board failed in the dome control.  A replacement board has been ordered.

TMO61 currently down.  The power company needs to do some repairs to the electric service for the observatory.  We don't know the schedule for finishing these repairs.

OC51 operational

We are working on porting, to other telescopes, the targets that were in the Coker30 queues. If you have an important object that needs to get moved quickly, please let Ken Menzies or I know, and we'll give it higher priority for porting.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484